OpenStudy (anonymous):

Simplify the expression: (1/1-cosx)+(1/1+cosx) -Please help! :D

OpenStudy (campbell_st):

why not 2 csc^2(x)

OpenStudy (anonymous):

Just like adding any other fractions together, your first step is to get a common denominator. This means that you multiply the first fraction by (1+cosx)/(1+cosx) and the second fraction by (1-cosx)/(1-cosx): 1*(1+cosx)/[(1-cosx)*(1+cosx)] + 1*(1-cosx)/[(1+cosx)*(1-cosx)] You can simplify the denominator by multiplying (using the FOIL method). It simplifies to 1- (cosx)^2. Simplifying the equation above you get: (1 + cosx + 1 - cosx)/[1 - (cosx)^2] The cosines in the numerator cancel each other out, leaving 2/[1 - (cosx)^2]. You can further simplify the denominator by using a trig identity: (sinx)^2 + (cosx)^2 = 1 Therefore, (sinx)^2 = 1 - (cosx)^2 By substituting (sinx)^2 for [1 - (cosx)^2] you get a final answer of 2/(sinx)^2

OpenStudy (anonymous):

the answer choices include: 2cscx csc^2x 2sec^2x 2csc^2x

OpenStudy (campbell_st):

well there you go, problem solved

OpenStudy (anonymous):

\[\huge \frac{1}{1-\cos x} + \frac{1}{1+\cos x}\] \[\huge \frac{1+\cos x +1-\cos x}{(1-\cos x)(1+\cos x)}\] \[\huge \frac{2}{1-\cos^2 x}=\huge \frac{2}{\sin^2 x} = 2cosec^2x\] @RavenclawUnited

OpenStudy (anonymous):

\[\huge 2cosec^2x = 2\csc^2 x\]

OpenStudy (anonymous):

how did you get to 2 csc^2x @campbell_st

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

\[\huge \frac{1}{sinx} = \csc x \rightarrow \frac{1}{\sin^2x} = \csc^2 x\] @RavenclawUnited

OpenStudy (campbell_st):

well its simply 2(1/sin(x))^2 and 1/sin = csc so substituting 2(csc(x))^2 which is written 2 csc^2(x)

OpenStudy (anonymous):

Thank you @campbell_st and @dpasingh